OpenStudy (anonymous):

Can someone explain how would I answer this question? Fans & Medals!! A line has a slope of and a y-intercept of 3.

OpenStudy (anonymous):

a slope of -1/2

OpenStudy (anonymous):

The equation of a line is represented by the equation y = mx + b, where "m" is the slope and "b" is the y-intercept. So y = -1/2 x + 3 if you just plug those values into the equation.

OpenStudy (anonymous):

Yea thats one way to do it. does anyone know how to do it another way?

OpenStudy (anonymous):

I just realized I don't know what the question is. What is the actual question asking?

OpenStudy (anonymous):

Just for the equation of the line?

OpenStudy (anonymous):

This should be in math, dear.

OpenStudy (anonymous):

Yes @fission-mailed and this is math @TheHyaenaQueen

OpenStudy (anonymous):

Not sure which specific way you are looking for then... There are a few different ways you can get the equation of the line, but y = mx + b is pretty standard and the easiest way in this case. You could also use the y - Y1 = m (x - X1) formula if it's point-slope form you want, but you usually learn point-slope form later on so I don't think that's what you want. Did you want it in standard form or something? Have you not learned y = mx + b yet?

OpenStudy (anonymous):

Yea I got the y=mx+b but its asking for more than one

OpenStudy (anonymous):

If you want it in standard form you can bring the y over and set the equation equal to 0. y = -1/2 x + 3 -1/2 x + 3 - y = 0 Rearrange because the y tends to go in the middle for standard form. -1/2 x - y + 3 = 0 If you want you can multiply by 2 to get rid of the fraction, which gives you the equation -x - 2y + 6 = 0

OpenStudy (anonymous):

The way I wrote it above is technically called general form, not standard, just an fyi for when it is set equal to 0 on one side (it's still a correct way to write the line). But if you want "Standard form" as in the name, only difference is it will just have the constant where the 0 is, so just move the 6 over the equal sign to get it in standard. -x - 2y = -6 You could divide the whole equation by -1 to make it look nicer, if you want. x + 2y = 6. There. Standard form. Whatever floats your bloat!